### v3

# Robust Multimodal Learning with Missing Modalities via Parameter-Efficient Adaptation

Abstract
Multimodal learning seeks to utilize data from multiple sources to improve the overall performance of downstream tasks. It is desirable for redundancies in the data to make multimodal systems robust to missing or corrupted observations in some correlated modalities. However, we observe that the performance of several existing multimodal networks significantly deteriorates if one or multiple modalities are absent at test time. To enable robustness to missing modalities, we propose a simple and parameter-efficient adaptation procedure for pretrained multimodal networks. In particular, we exploit modulation of intermediate features to compensate for the missing modalities. We demonstrate that such adaptation can partially bridge performance drop due to missing modalities and outperform independent, dedicated networks trained for the available modality combinations in some cases. The proposed adaptation requires extremely small number of parameters (e.g., fewer than 0.7% of the total parameters) and applicable to a wide range of modality combinations and tasks. We conduct a series of experiments to highlight the missing modality robustness of our proposed method on 5 different datasets for multimodal semantic segmentation, multimodal material segmentation, and multimodal sentiment analysis tasks. Our proposed method demonstrates versatility across various tasks and datasets, and outperforms existing methods for robust multimodal learning with missing modalities.
1 Introduction
Multimodal learning (MML) [4, 63] is a general framework for processing, combining, and understanding information from multiple, diverse data sources. Fusing knowledge from multiple modalities (e.g., text, images, audio, and sensor data) is expected to provide more accurate and reliable systems. In recent years, MML has achieved remarkable success in a wide range of applications, including image segmentation [6, 55, 70], captioning [71, 65], classification [15, 46], sentiment analysis [50, 28], and autonomous driving [61, 45]. In all these applications, one often encounters situations where some modalities are corrupted or missing due to hardware limitations/failures, privacy concerns or data acquisition cost/constraints. The ability to handle corrupt or missing modalities is thus crucial for the robustness and reliability of multimodal systems. However, most of the existing multimodal models are not designed to handle corrupt or missing modalities. The primary focus of this paper is to study and enhance robustness of existing multimodal models in different missing modality scenarios.
Recent studies [38, 18, 41] have shown that MML is not inherently robust to missing modalities and performance can drop significantly when modalities are missing at test time. Existing approaches for robust MML usually work for specific combinations of modalities they are trained for and tend to perform poorly when applied to untrained combinations. For instance, one approach is to adopt robust training strategies such as modality dropout during training [42, 25], partial or full modality masking [2, 48], and knowledge distillation [52, 40]. These approaches either require specialized training strategies or utilize extra models/sub-networks to guide the underlying model. Another approach replaces uninformative tokens with aggregated informative tokens from different modalities or learns to predict tokens for the specific missing modalities [55, 58, 48]. Training such separate (independent) networks for every possible modality combination is not feasible specially when the number of input modalities is large. One recent approach for robust MML is to impute missing modalities from the available modalities [64, 47, 11]. Performance of these methods depend on the generation model that imputes the missing modalities.
In this paper, we propose a parameter-efficient approach to adapt existing multimodal networks to perform well on different missing modality scenarios. Our main objective is to modify the network in a controllable manner as a function of available modalities. For instance, if a modality is missing, we seek to modify how the features from available modalities are extracted and fused for the inference. Instead of learning an independent network for each modality combination, our goal is to perform parameter-efficient adaptation. Figure 1 illustrates our proposed method, where a given multimodal network can be adapted to arbitrary modality combinations by transforming the intermediate features of the available input modalities at different layers. To achieve parameter-efficient adaptation, we propose to use simple linear transformations such as scaling and shifting to modulate the intermediate features or low-rank increments of features to compensate for the missing modalities. Our method does not require retraining the entire model or any specialized training strategy. The adapted networks provide significant performance improvement over the multimodal networks trained with all modalities and tested with missing modalities. Performance of the adapted models is also comparable or better than the models that are exclusively trained for each input modality combination. We present a series of experiments to evaluate our method and compare with existing methods for robust MML. We tested different parameter-efficient adaptation strategies and found intermediate feature modulation with scaling and shifting provides overall best performance with less than 0.7% of the total parameters.
Contributions. The main contributions can be summarized as follows.
  • We propose parameter-efficient adaptation procedure for multimodal learning that is robust to missing modalities. The adapted model can easily switch to different network states based on the available modalities with minimal latency, computational, or memory overhead.
  • The adapted networks provide notably improved performance with missing modalities when compared to models trained with all modalities and is comparable to or better than the networks trained for specific modality combinations.
  • Our approach is versatile and adaptable to a wide range of multimodal tasks and models. Detailed evaluations on different datasets and tasks show that our approach outperforms existing baseline methods and robust models designed for specific tasks and datasets.
Figure 1: a) Overview of our approach for robust multimodal learning with missing modalities via parameter-efficient adaptation. A model pretrained on all the modalities is adapted using a small number of learnable parameters to handle different modality combinations. We insert adaptable layers after each layer of the encoders and the fusion block to learn the modulation as a function of the available input modalities to compensate for the missing modalities. The grayed-out branch (missing modality branch) is inactive and does not contribute to the output. b) Low-rank model adaption computes features using frozen weights and low-rank weight updates and combine them. c) Scale and shift feature adaptation transforms input by element-wise multiplication and addition. One set of parameters is learned for each modality combination.
2 Related Work
Multimodal learning with missing modalities has been studied for different applications in recent years. For instance, robustness in vision-language tasks with multimodal transformers in [38], multimodal sentiment analysis in [41], multimodal classification in [18], and multimodal action recognition in [58]. These studies have shown that the task performance can drop significantly when modalities are missing during test time.
Robust training strategies have been proposed to make models robust to different missing modalities. Such approaches include modality dropout during training [42, 25], unified representation learning [29], and supervised contrastive learning [14]. Modality masking during training has become a popular choice for enhancing robustness. [48] utilized complementary random masking, [21] used masked auto encoder, and [38] applied masked cross-modal attention for enhancing robustness of the underlying model. [18] proposed noisy perturbation of modalities during training for robust multimodal sentiment analysis. Recently, [30] proposed uni-modal ensemble with modality drop and substitution augmentation during training to adapt to different missing modality scenarios.
Design of robust models and fusion strategies is another approach for robust MML. [13] proposed a recursive meshing technique called SpiderMesh and [48] designed complementary random masking (CRM) and knowledge distillation based framework for robust RGB-thermal semantic segmentation. [55] proposed TokenFusion to dynamically detect and replace uninformative tokens with projected tokens from other modalities for robust RGB-depth semantic segmentation, image-to-image translation, and 3D object detection. [54] proposed a model that learns modality-shared and modality-specific features for robust brain tumour segmentation. [7] proposed a robust fusion strategy for multimodal classification. The main limitation of these methods is that they are generally designed for a specific modality combination and do not perform well when applied to other multimodal tasks [33].
Knowledge distillation and generation methods have also become popular for robust MML. Studies by [47] and [64] used GAN based generative models while [11] used VAE based generative models for imputing missing modalities from available input modalities for unlerlying multimodal tasks. Recently [44] introduced an approach to learn missing modality tokens from available modalities. Different knowledge distillation approaches have also been applied in several multimodal tasks. [52] proposed mean teacher and [40] introduced multimodal teacher for semi-supervised image segmentation. [48] and [18] applied self-distillation loss for robust RGB-thermal semantic segmentation. Apart from these approaches, weight space ensembling [59], policy learning [38] and optimal fusion strategy designing [40] were also studied for robust MML for various tasks.
These approaches are either designed for specific tasks/modality combinations [55, 48, 13] or require training extra modules/sub-networks [52, 40] for guiding the model under different missing modality scenarios. Our goal is to design a generic framework that is parameter-efficient and applicable to any model and modality combinations.
Parameter-efficient network adaptation has become very popular for uni-modal network adaptation in recent years [23, 34]. A number of parameter-efficient methods have been proposed for transfer learning [20, 10] and domain/task adaptation [24, 31]. We can divide the approaches into two major categories: low-rank/additive approach and feature transformation approach.
Low-rank/additive adaptation has been applied for uni-modal model fine-tuning and domain adaptation. For instance, LoRA [24], QLoRA [9], KronA [12] and KAdaptataion [22] learn low-rank factors for task/domain adaptation. Let $W$ be the weight matrix of any dense layer of a given pretrained uni-modal model. These approaches learn a low-rank weight update matrix $\Delta W$ to transform the input $x$ to that layer as $h=Wx+\Delta Wx$, where $h$ is the updated feature. Since the update matrix $\Delta W$ is low-rank, the number of learnable parameters remains a fraction of the total number of model parameters. BitFit [5] is another approach that adapts the bias terms while keeping everything else frozen.
Feature modulation based approach is another parameter-efficient method to transform intermediate features of the pretrained model [26, 1, 60, 31]. As shown in Figure 1c, it applies a linear transformation to the given input token/feature with learnable scale ($\gamma$) and shift ($\beta$) parameters. Given an input token $x$, this approach generates the output token as $h=\gamma\odot x+\beta$, where $\gamma,\beta,x,h$ are vectors of same dimension and $\odot$ represents element-wise multiplication along the embedding dimension. These scale ($\gamma$) and shift ($\beta$) parameters are input-independent and learned during the training process to help the model adjust and fine-tune its representations for better performance on the underlying task.
Though parameter-efficient adaptation approaches have shown great potential in transfer learning, model fine tuning and task/domain adaptation, their potential remains unexplored in the context of missing modality in MML. In this study, we focus on parameter-efficient approaches to build our generic framework to enhance missing modality robustness in MML.
3 Proposed Method
In this section we first present a general framework for network adaptation for missing modalities. Then we discuss why we focus on parameter-efficient adaptation, present details of our proposed approach for missing modality adaptation and highlight the key benefits of our approach.
3.1 Network Adaptation for Missing Modalities
Let us denote the set of input modalities for a given multimodal task as =$\mathcal{M}=\{m_{1},\ldots,m_{M}\}$. Given the full set $\mathcal{M}$, we can train a model $f$ with parameters $\Theta_{\mathcal{M}}$ that maps inputs from all the modalities (denoted as $\mathcal{X_{M}}$) to an output $y_{\mathcal{M}}$ as
$y_{\mathcal{M}}=f(\mathcal{X_{M}};\Theta_{\mathcal{M}}).$ (1)
While we can ensure the availability of all input modalities during training, it is possible that some modalities may be inaccessible at test time, especially after real-world deployment. Any subset of modalities $\mathcal{M}$ can get missing due to hardware failure, data acquisition cost or privacy concerns. If we use a model trained on all the input modalities as denoted by (1), significant performance drop is observed when a subset of modalities gets missing during test time as shown in Table 1.
3.1.1 Naïve approach
When a subset of the modalities $\mathcal{M}$ is missing, a simple and naïve approach is to train a new model for the available input modalities. Without loss of generality, suppose $\mathcal{K}\subset\mathcal{M}$ represents missing modalities. We can use the available input modalities $\mathcal{S}=\mathcal{M}\setminus\mathcal{K}$ to retrain the model $f$ for a new set of parameters $\Theta_{\mathcal{S}}$ as
$y_{\mathcal{S}}=f(\mathcal{X_{S}};\Theta_{\mathcal{S}}),$ (2)
where $\mathcal{X_{S}}$ represents input data for modalities in $\mathcal{S}$. In principle, we can train one model for every possible $\mathcal{S}\subset\mathcal{M}$ and use the corresponding model at the test time. Such an approach is infeasible because of computational and storage resources required to train models for a large number of possible modality combinations. Furthermore, deploying a large number of trained models and selecting one of them at test time is not feasible in real-world scenarios. Another drawback of this method is that, even though we would like $y_{\mathcal{S}}\approx y_{\mathcal{M}}$, the training process mentioned earlier does not guarantee it.
3.1.2 Parameter-efficient approach
We propose an alternative approach to adapt a single model for all subsets of input modalities $\mathcal{S}\subset\mathcal{M}$ in a parameter-efficient manner. First, we select a model $f$ trained on the full set of modalities $\mathcal{M}$ as shown in (1) and freeze the parameters $\Theta_{\mathcal{M}}$. Then we learn a small number of parameters $\Delta_{\mathcal{S}}$, specific to the available input modality set $\mathcal{S}$, and update the model as
$\hat{y}_{\mathcal{S}}=f(\mathcal{X_{S}};\Theta_{\mathcal{M}},\Delta_{\mathcal{S}}),$ (3)
where $\hat{y}_{\mathcal{S}}$ represents the prediction of the updated model. Our goal is to keep $\hat{y}_{\mathcal{S}}$ close to all modality prediction ${y}_{\mathcal{M}}$ in the best case ($\hat{y}_{\mathcal{S}}\approx{y}_{\mathcal{M}}$) and close to the prediction ${y}_{\mathcal{S}}$ made by a model trained on the available input modalities in the worst case ($\hat{y}_{\mathcal{S}}\approx{y}_{\mathcal{S}}$).
The adaptation method shown in (3) is considered parameter-efficient if the number of parameters in $\Delta_{\mathcal{S}}$ is significantly smaller compared to the total number of parameters in $\Theta_{\mathcal{M}}$. During adaptation, we only train the learnable parameters $\Delta_{\mathcal{S}}$. In our experiments, as we discuss later, we keep $\Theta_{\mathcal{M}}$ frozen and demonstrate that less than $0.7\%$ of the total parameters for $\Delta_{\mathcal{S}}$ are sufficient for network adaptation (see supplementary material for details).
3.1.3 Need for parameter-efficient adaptation
In the last few years, a number of approaches have been proposed for MML with missing modalities. To the best of our knowledge, parameter-efficient adaptation is still unexplored in this field. The current methods for robust MML, as discussed in Section 2, require retraining the whole model with specialized training strategy [48, 25] or utilize extra module/sub-network to guide the multi-modal model [40, 52]. Furthermore, these methods are not very generic and do not perform well on different missing modality scenarios as shown in Table 2 and 3. To solve these issues, we propose parameter-efficient adaptation for enhancing robustness of MML in the face of missing modalities. Our approach requires learning a very small number of parameters for different missing modality scenarios without the need to retrain the whole network. Furthermore, it is also applicable to diverse model architectures, tasks and modality combinations.
3.2 Parameter-Efficient Adaptation for Missing Modalities
In this section we first discuss how we adapt a given multimodal network for different missing modality scenarios. Then we discuss the reasons behind selecting intermediate feature modulation as the adaptation technique. Finally, we compare it with other parameter-efficient adaptation methods and discuss about some key benefits of our selected approach.
Adaptation for multimodal models. To the best of our knowledge, no parameter-efficient adaptation approach has been proposed or applied for multimodal model adaptation to handle missing modalities. We draw our motivation from low-rank adaptation and feature modulation based approach to compensate for the missing modalities. For instance, LoRA [24], BitFit [5], layer/feature normalization [1, 26], and scale and shift features (SSF) [31] approaches can enhance the representation capabilities of deep models for domain adaptation and model fine-tuning. We extend these adaptation approaches to build a generic framework that can transform the intermediate features of the available modalities to find an optimal feature representation to compensate for the performance gap due to missing modalities.
3.2.1 Training: model adaptation for missing modalities
Our approach for multi-modal model adaptation for missing modality is illustrated in Figure 1. Without loss of generality, let us assume a generic multi-modal model in which each modality goes through a separate modality-specific encoder. The extracted features from all the modalities are fused with a fusion block and the fused feature is passed to a decoder head for making prediction. Each encoder and the fusion block have multiple layers (linear/convolutional/norm layers). This setup can be easily generalized to models with shared encoder and/or different (early or mid) fusion strategy. We also do not make any assumption about the underlying model architecture.
We train this multimodal network $f$ with all available modalities in $\mathcal{M}$ to learn the parameters $\Theta_{\mathcal{M}}$ as shown in (1). Then we adapt $f$ for different subsets of available modalities $\mathcal{S}\subset\mathcal{M}$. Unlike existing methods, we do not try to generate [39, 2], approximate [55, 54] or distill knowledge [48, 40] from any other modality/sub-network. Our goal is to learn a modified function for the available input modalities to appropriately learn and fuse features to compensate for any missing modality. Instead of re-training the entire network on the available modalities as shown in (2), we adapt the base network $f$ and focus on learning a minimal set of parameters following (3).
[…]
Below we discuss how to use low-rank and intermediate feature modulation-based multimodal network adaptation to accommodate missing modalities. We believe our framework is generic and can also incorporate other parameter-efficient adaptation approaches.
Low-rank/additive adaptation. We extend low-rank/additive approaches to adapt multimodal model for missing modalities. Let us assume that $W_{m}$ be one of the weight matrices from any layer for the $m^{th}$ input modality where $m\in\mathcal{S}$. As shown in Figure 1b, we learn a low-rank weight update matrix $\Delta W_{m}$ for that layer to transform the input $h_{m,i}$ to the layer as
$h_{m,o}=W_{m}h_{m,i}+\Delta W_{m}h_{m,i},\text{ for all }m\in\mathcal{S},$ (4)
where $h_{m,o}$ is the transformed output feature that is passed to the next layer in the model. Since the weight update matrix $\Delta W_{m}$ is low-rank, the total number of learnable parameters remains a fraction of the total number of model parameters. We can represent the learnable parameters $\Delta_{\mathcal{S}}=\{\Delta W_{m}\}_{m\in\mathcal{S}}$ as the collection of all low-rank update matrices.
Intermediate feature modulation. We extend SSF [31] method to work with multimodal models with missing modalities. The adaptable SSF layers modulate the intermediate tokens/features from each available modality at every layer as shown in Figure 1c. For the $m^{th}$ input modality where $m\in\mathcal{S}$, we denote the learnable scale parameters as $\gamma_{m}\in\mathbb{R}^{d}$ and the learnable shift parameters as $\beta_{m}\in\mathbb{R}^{d}$ where $d$ is the embedding dimension of the model. Since we insert SSF layers after each frozen layer in the pretrained model $f$, the output ${h}_{m,o}\in\mathbb{R}^{N\times d}$ form any frozen layer for the $m^{th}$ input modality goes through the SSF layer that follows it. The SSF layer applies a linear transformation on ${h}_{m,o}$. We can formulate the linear transformation for multimodal input as follows:
${h}_{m,i}=\gamma_{m}\odot{h}_{m,o}+\beta_{m},\text{ for all }m\in\mathcal{S},$ (5)
where ${h}_{m,i}\in\mathbb{R}^{N\times d}$ is the transformed feature which is fed to the next frozen layer in the model and $N$ is the number of tokens. Note that if the output of any layer is of shape $(H\times W\times d)$ (for convolutional layers), we reshape it to $(N\times d)$, where $N=H\times W$, before applying the linear transformation. We reshape the transformed feature back to the original shape (if required) before passing it to the next layer. We can represent the learnable parameters as $\Delta_{\mathcal{S}}=\{\gamma_{\mathcal{S}},\beta_{\mathcal{S}}\}=\{\gamma_{m},\beta_{m}\}_{m\in\mathcal{S}}$. BitFit [5] method can also be used for adaptation as we only need to learn the bias/shift terms $\beta_{m}$ for all $m\in\mathcal{S}$. We modify (5) as
${h}_{m,i}={h}_{m,o}+\beta_{m},\text{ for all }m\in\mathcal{S},$ (6)
and the learnable parameters can be represented as $\Delta_{\mathcal{S}}=\{\beta_{m}\}_{m\in\mathcal{S}}$. Thus the intermediate features from each available modality are modulated to find a better representation to compensate for the missing modalities.
3.2.2 Inference: model adaptation for missing modalities
At the test time, we load the base multimodal model $f$ with the pretrained weights $\Theta_{\mathcal{M}}$. If all the modalities are available, then we can use $\Theta_{\mathcal{M}}$ to make predictions. When a subset of the modalities are missing, we can select the learned parameters $\Delta_{\mathcal{S}}$ corresponding to the available input modalities $\mathcal{S}$, insert them into the model and use them to make prediction as follows:
$\hat{y}_{\mathcal{S}}=\begin{cases}f(\mathcal{X_{S}};\Theta_{\mathcal{M}})&\text{if }\mathcal{S}=\mathcal{M},\\ f(\mathcal{X_{S}};\Theta_{\mathcal{M}},\Delta_{\mathcal{S}})&\text{if }\mathcal{S}\subset\mathcal{M}.\end{cases}$ (7)
Since we are inserting the adaptable layers after each layer, it does not require any major change to the model architecture and can be done easily without reloading all the model parameters $\Theta_{\mathcal{M}}$. We just need to load the parameters in $\Delta_{\mathcal{S}}$ and insert them into the model. Since we only insert a very small number of additional parameters (less than 0.7% of the total model parameters), it adds very limited computational overhead. Furthermore, if a different subset of modalities becomes available, the adjustment is straightforward. We only need to replace the existing learned parameters $\Delta_{\mathcal{S}}$ with the corresponding parameters for the available modality set, ensuring adaptability and flexibility in handling diverse combinations of available modalities during the testing phase.
We only insert adaptable layers in the encoders and fusion blocks, while keeping the decoder/prediction head unchanged. We observed that using pretrained decoder/prediction head provided a good overall performance with several missing modalities.
3.2.3 Feature modulation vs low-rank adaptation
While we present three adaptation approaches in (4), (5), and (6), we select intermediate feature modulation with SSF (5) as the main approach for our experiments. We primarily selected this technique for robust MML with missing modalities because of its simplicity and effectiveness. Our experiments show that feature transformation via simple linear transformation with SSF works well for most of the scenarios compared to other parameter-efficient adaptation approaches as summarized in Table 4. We provided a detailed comparison in terms of mean accuracy, F1 score and % mIoU in Table 8, 9 and 10 in the supplementary section. SSF shows great promise in enhancing representation power [1], faster convergence [26], prevents loss of information in the representation learning process [60] and mitigates distribution mismatch between the upstream and downstream tasks [31]. These characteristics motivated us to extend this method for multimodal learning with missing modalities and build a generic framework that is very effective in learning the proper modulation of available input modalities to bridge the performance gap in the face of missing modalities.
Some key benefits of this approach are as follows. First, The parameters $\gamma,\beta$} are independent of the input features/modalities, which makes it applicable to diverse tasks and input modality combinations. Second, we can easily insert these learnable layers in existing models without changing the model architecture. We can easily switch/select the corresponding SSF parameters for a given input modality combination. Finally, it introduces extremely small number of additional learnable parameters. The resulting adaptation offers significant savings compared to training a separate model for each input combination or retraining the model using some specialized training strategy like modality dropout [25, 42] or knowledge distillation [48, 40].
4 Experiments and Results
We performed detailed experiments to evaluate the performance of our proposed method for multimodal semantic segmentation, multimodal material segmentation and multimodal sentiment analysis tasks on five datasets. We also present comparison with existing baseline methods that are robust to missing modalities.
Table 1: Performance of Pretrained, Dedicated, and Adapted networks with missing modalities. CMNeXt is the base model for multimodal semantic segmentation for MFNet and NYUDv2 datasets and multimodal material segmentation for MCubeS dataset. HHA-encoded images were used instead of raw depth maps. Bold letters represent best results.
Dataset Input Missing Pretrained Dedicated Adapted (Ours)
MFNet RGB-Thermal - 60.10 60.10 -
RGB Thermal 53.71 55.86 55.22
Thermal RGB 35.48 53.34 50.89
NYUDv2 RGB-Depth - 56.30 56.30 -
RGB Depth 51.19 52.18 52.82
Depth RGB 5.26 33.49 36.72
MCubeS RGB-AoLP-DoLP-NIR - 51.54 51.54 -
RGB-AoLP-DoLP NIR 49.06 49.48 51.11
RGB-AoLP DoLP-NIR 48.81 48.42 50.62
RGB AoLP-DoLP-NIR 42.32 48.16 50.43
4.1 Datasets
Multimodal semantic segmentation. We experimented on two datasets for multimodal semantic segmentation tasks: MFNet dataset [17] with RGB-Thermal input modalities and NYUDv2 dataset [49] with RGB-Depth input modalities. These datasets are divided into train and test sets along with ground truth per-pixel annotation for the underlying semantic segmentation tasks.
Multimodal material segmentation. We used MCubeS dataset [32] for multimodal material segmentation having RGB, Angle of Linear Polarization (AoLP), Degree of Linear Polarization (DoLP) and Near-Infrared (NIR) images as input modalities. The dataset is divided into train, validation and test sets along with ground truth per-pixel annotation for 20 material classes.
Multimodal sentiment analysis. CMU-MOSI from [69] and CMU-MOSEI from [3] are two popular datasets for multimodal sentiment analysis using audio, visual and text as input modalities. They contain 2199 and 23453 annotated data samples, respectively, divided into train, validation, and test sets.
Figure 2: Examples of predicted segmentation maps for the Pretrained and Adapted models for multimodal semantic segmentation (on MFNet and NYUDv2) and material segmentation (on MCubeS). Title above each subimage shows method name (tested with available modalities). CMNeXt column shows the predictions with all the modalities. Segmentation quality improves significantly after model adaptation for all the input modality combinations. Green boxes highlight areas with salient differences in results (e.g., cars and humans missing in the Pretrained model with missing modalities but visible in the Adapted model). For MCubeS dataset, we only show RGB images and A, D and N denote angle of linear polarization, degree of linear polarization, and near-infrared, respectively.
4.2 Implementation Details
To investigate missing modality adaptation performance in multimodal semantic and material segmentation tasks, we use the CMNeXt [70] as the base model. We use multimodal transformer [53] as the base model for multimodal sentiment analysis. We train the corresponding base model with all the input modalities for each dataset. To evaluate performance with missing modalities, we provide the available modalities and set the missing modalities to zero. To perform model adaptation for any modality subset $\mathcal{S}\subset\mathcal{M}$, we freeze the pretrained weights and insert learnable SSF layers. Then we fine tune the learnable parameters for 100 epochs for multimodal segmentation tasks and until convergence for multimodal sentiment analysis tasks.
For multimodal segmentation tasks, we set the initial learning rate to $6\times 10^{-5}$ and applied a polynomial learning rate scheduler with a power of 0.9. The first 10 epochs were set as the warm-up, during which the learning rate was set to 0.1 times the original rate. The scale parameters $(\gamma)$ were initialized with all 1s and the shift parameters $(\beta)$ were initialized with all 0s. We used cross-entropy loss function and AdamW optimizer as proposed in [37], with an epsilon value of $10^{-8}$ and a weight decay of 0.01. We used a batch size of 4 and report single scale performance for all the datasets. All other hyper-parameters and configurations are the same as [70]. For multimodal sentiment analysis tasks, we used the default settings for the datasets as configured in the codebase [66]. We have included additional details for each dataset and experimental setup in the supplementary section.
4.3 Baseline Methods
We report experiments and results for different methods that are listed as follows. Pretrained model refers to the base model that is trained with all the available modalities. We provide zeros for the missing modalities. Modality Duplication means that one of the available modalities is used as a substitution for the missing modality. Dedicated training refers to independent models trained for each input modality combination. Adapted model refers to the model that is adapted using our approach for each input modality combination.
Different robust methods have been proposed for different multimodal tasks. We compare our method with the following methods: masking and recursive meshing based approach SpiderMesh [13], variational probabilistic fusion based approach VPFNet [33], modality discrepancy reduction approach MDRNet [72], knowledge distillation and robust training based approach CRM [48] for robust RGB-thermal semantic segmentation. Dynamic channel exchange based approach CEN [56], dynamic token selection/generation based approach TokenFusion [55], assymetric fusion based approach AsymFusion [57] for robust RGB-depth semantic segmentation.
For every task/dataset, we adopted the experimental setup used in the corresponding previous studies. We used the reported results from prior works where possible. It is important to note that, because of this criteria, some of the baseline methods may only be present in specific experiments depending on the availability of their reported numbers. We also perform detailed comparison of SSF with other parameter-efficient adaptation techniques. For all the experiments we follow the same setup suggested by the corresponding papers and report mean accuracy (mAcc), F1 score and mean intersection over union (mIoU) when available.
Ablation studies and comparison of different parameter-efficient adaptation methods show that SSF-based adaptation provides overall best performance. We summarize the results for scale only, shift only, BitFit [5], norm layer fine-tuning and LoRA [24] in Table 4. We also add detailed comparison for all the datasets in the supplementary section of the paper.
4.4 Experiments for Multimodal Segmentation
In this section, we present experiments for multimodal semantic segmentation on MFNet and NYUDv2 datasets, and multimodal material segmentation on MCubeS dataset. First, we show an overall comparison of our approach with existing baselines on Table 1 and then we compare our approach with existing robust methods on Table 2 and 3. We report the detailed results for multimodal material segmentation along with per class % intersection over union (IoU) comparisons between the Pretrained and Adapted models in the supplementary section of the paper.
4.4.1 Overall performance comparison
We present the performance comparison of Pretrained, Modality Duplication, Dedicated, and Adapted networks for different missing modalities in Table 1. We observe that the performance of the Pretrained model drops significantly with missing modalities. We see a 6.39% drop when Thermal is missing on MFNet dataset and 5.11% drop when Depth is missing on NYUDv2 dataset compared to the case when all modalities are available. The effect is amplified when RGB gets missing as we observe 24.62% drop on MFNet dataset and 51.04% drop on NYUDv2 dataset. On MCubeS dataset, we observe 2.48–9.22% drop in pretrained model when different modality combinations are missing. Similar trend of performance drop is observed for modality duplication approach though it performs better than pretrained models for most of the cases.
The overall performance of Adapted models with missing modalities is significantly better than Pretrained models and Modality Duplication approach. For MFNet, an improvement of 1.51% compared to the Pretrained model is observed when RGB is available and thermal is missing. The Adapted model performance is also close to the performance of Dedicated network trained for RGB only. The adapted model shows a significant improvement of 15.41% compared to the Pretrained model when RGB is missing. For NYUDv2 dataset, we see 1.63% and 31.46% performance improvement compared to Pretrained model when depth and RGB are missing, respectively. In both cases, the performance of the Adapted model is better than the Dedicated model. For all input combinations on MCubeS dataset, we see 1.82–8.11% performance improvement compared to the Pretrained model. The Adapted model performs better than Dedicated models in all the cases.
Table 2: Performance comparison with existing robust methods for MFNet dataset. RGB and Thermal columns report performance when only RGB and Thermal are available. Average column reports average performance when one of the two modalities gets missing. ‘-’ indicates that results for those cells are not published. Mean accuracy (mAcc) and % mean intersection over union (mIoU) are shown for all the experiments.
Methods RGB Thermal Average
mAcc % mIoU mAcc % mIoU mAcc % mIoU
FuseNet [19] 11.11 10.31 41.33 36.85 26.22 23.58
MFNet [17] 26.62 24.78 19.65 16.64 23.14 20.71
RTFNet [51] 44.89 37.30 26.41 24.57 35.65 30.94
SAGate [6] 32.01 30.57 13.34 12.51 22.68 21.54
FEANet [8] 15.96 8.69 58.35 48.72 37.16 28.71
MDRNet [72] 57.11 45.89 41.98 30.19 49.55 38.04
VPFNet [33] 48.14 41.08 42.20 35.80 45.17 38.44
SpiderMesh [13] - 39.60 - 50.50 - 45.05
CRM [48] - 52.70 - 53.10 - 52.90
Adapted (Ours) 67.18 55.22 66.70 50.89 66.94 53.06
Table 3: Performance comparison with existing robust methods for NYUDv2 dataset. RGB and Depth columns report performance when only RGB and Depth are available. Average column indicates average performance when one of the two modalities gets missing. $\ast$ indicates that available codes and pretrained models from the authors were used to generate the results. Other results are from the corresponding papers.
Methods RGB Depth Average
mAcc % mIoU mAcc % mIoU mAcc % mIoU
FCN [36] 44.70 31.60 35.70 25.20 40.20 28.40
Dilated FCN-2s [27] 47.10 32.30 39.30 26.80 43.20 29.55
AsymFusion (R-101) [57] 59.00 46.50 45.60 34.30 52.30 40.40
CEN (R-101) [56] $\ast$ 51.77 39.59 28.98 19.32 40.38 29.46
TokenFusion [55] $\ast$ 63.49 49.32 46.83 36.84 55.16 43.08
Adapted (Ours) 67.96 52.82 52.42 36.72 60.19 44.77
4.4.2 Performance comparison for RGB-Thermal semantic segmentation on MFNet dataset
We compare the performance of the Adapted model with existing robust models for RGB-thermal semantic segmentation on MFNet dataset in Table 2. The results show that the Adapted model offers the best average performance compared to the other baseline models, in terms of mean accuracy and % mIoU. Among the robust models, robust training and knowledge distillation based model CRM [48] shows competitive performance with the Adapted model. The Adapted model performs better when only RGB is available while CRM performs better when only Thermal is available. Notably CRM is designed specifically for RGB-Thermal pairs and requires specialized training approach and utilizes self-distillation loss between the clean and masked modalities to train the model. In contrast, our approach is applicable to any input modalities and does not require any specialized training technique. Our approach performs significantly better compared to masking and recursive meshing based SpiderMesh [13], variational pobabilistic fusion based VPFNet [33] and modality discrepancy reduction based MDRNet [72] models.
4.4.3 Performance comparison for RGB-Depth semantic segmentation on NYUDv2 dataset
Table 3 shows the performance comparison with existing robust models for RGB-Depth semantic segmentation on NYUDv2 dataset. For our experiments, we use HHA-encoded images proposed by [16] instead of raw depth maps. On an average, the Adapted model performs better than the existing robust models. Dynamic token selection and substitution based model TokenFusion [55] performs slightly better (+0.12%) in terms of mIoU when Depth is available and RGB is missing, but shows larger drop (-5.59%) in mean accuracy. On the other hand, the Adapted model performs significantly better (+3.5% mIoU and +4.47% mean accuracy) when RGB is avalable and Depth is missing. The average performance of the Adapted model is also better than the TokenFusion model despite the fact that TokenFusion was designed to work with RGB-Depth pair, whereas the Adapted method is independent of input modalities. Our method also performs significantly better compared to dynamic channel exchange based CEN [56] and asymmetric fusion based AsymFusion [57] models.
4.4.4 Visualization of predictions
For qualitative analysis, we show some examples of the predicted segmentation maps form the Pretrained and Adapted models in Figure 2. For each dataset, we show the input images, predictions when all the modalities are available during both training and test time (CMNeXt column), predictions from the pretrained and adapted models for different available/missing modality scenarios (Available input modality names are shown in parentheses above each image). We see in Figure LABEL:fig:vis-mfnet-1, the Pretrained model fails to detect humans when only RGB images are available and cars when only Thermal images are available. The adapted model can detect both humans and cars with missing modalities.
For NYUDv2 dataset, as shown in Figure LABEL:fig:vis-nyu-1, the Adapted model can detect window, bed, and furniture with higher accuracy than the Pretrained model with missing modalities. We only show the RGB input images for MCubeS dataset for brevity in Figure LABEL:fig:vis-mcubes-1. The Adapted model can identify sand, sky, and gravel with higher accuracy than the pretrained model. In all cases, the predictions from the Adapted model with missing modalities are closer to the predictions of the pretrained model with all the input modalities. We provide additional visualizations in the supplementary section of the paper.
Table 4: Performance comparison (mIoU) of different parameter-efficient adaptation techniques for MFNet, NYUDv2, and MCubeS datasets. Each column reports mIoU of the Adapted model with the corresponding modalities, and Avg indicates average performance. A and D denote Angle and Degree of Linear Polarization.
Datasets MFNet NYUDv2 MCubeS
Methods RGB Thermal Avg RGB Depth Avg RGB RGB-A RGB-A-D Avg
Pretrained 53.71 35.48 44.60 51.19 5.26 28.23 42.32 48.81 49.06 46.73
Dedicated 55.86 53.34 54.60 52.18 33.49 42.84 48.16 48.42 49.48 48.69
Scale Only 54.77 49.23 52.00 53.04 36.12 44.58 50.16 50.55 51.13 50.61
Shift Only 54.57 48.96 51.77 53.04 36.25 44.65 50.13 50.40 50.86 50.46
BitFit 54.39 49.07 51.73 53.09 36.64 44.87 50.19 50.57 51.07 50.61
LoRA 54.19 47.45 50.82 52.87 34.97 43.92 49.59 50.07 50.80 50.15
Norm 54.65 47.49 51.07 53.05 34.73 43.49 49.95 50.51 51.07 50.51
Scale and Shift 55.22 50.89 53.06 52.82 36.72 44.77 50.43 50.62 51.11 50.72
Table 5: Comparison of our adaptation technique with existing methods for multimodal sentiment analysis.
Datasets Methods Audio Visual Audio-Visual Average
ACC F1 ACC F1 ACC F1 ACC F1
CMU -MOSI MulT ([53]) 48.31 40.98 52.44 51.77 48.93 41.95 49.89 44.90
MFN ([68]) 56.86 44.81 55.95 42.94 56.86 51.07 56.56 46.27
TFN ([67]) 42.23 25.07 42.38 25.40 42.23 25.07 42.28 25.18
BERT_MAG ([43]) 57.77 42.31 57.77 42.31 57.77 42.31 57.77 42.31
LMF ([35]) 42.23 25.07 43.14 27.54 43.29 27.61 42.89 26.74
Adapted (Ours) 50.00 46.71 54.88 54.39 55.49 53.96 53.46 51.69
CMU - MOSEI MulT ([53]) 37.15 20.12 38.28 23.70 41.91 32.78 39.11 25.53
MFN ([68]) 58.48 58.31 60.35 59.48 59.74 60.37 59.52 59.39
TFN ([67]) 37.15 20.12 37.15 20.12 37.15 20.12 37.15 20.12
BERT_MAG ([43]) 62.83 48.50 61.39 49.70 62.83 48.51 62.35 48.90
LMF ([35]) 42.38 34.48 57.15 57.85 55.94 56.63 51.82 49.65
Adapted (Ours) 62.85 55.55 62.49 60.00 63.32 60.69 62.89 58.75
4.4.5 Performance comparison with parameter-efficient model adaption techniques
Apart from robust models, we also compare different parameter-efficient adaptation techniques. We summarize the results in Table 4. For MFNet dataset, SSF outperforms all the methods and performance is significantly better than the Pretrained model and close to the Dedicated models. For NYUDv2 and MCubeS datasets, the Adapted model performs better than both Pretrained and Dedicated models. These experiments also show that SSF performs better than other methods for most of the input modality combinations for all the datasets. We show a detailed comparison for each dataset in terms of mean accuracy, F1 score and % mIoU in the supplementary section of this paper.
4.5 Experiments for Multimodal Sentiment Analysis
We tested our adaptation method for multimodal sentiment analysis on CMU-MOSI [69] and CMU-MOSEI [3] datasets, and report the results in Table 5. We use multimodal transformer (MulT) [53] as the base model and adapt it using our approach. We observed that when text is available and either audio or video or both are missing at the test time, the performance does not drop significantly. Similar trend was reported in [18]. If text is missing at test time, then the performance of the Pretrained model drops significantly. The Adapted models can partially compensate for missing text and offer significantly better performance compared to the pretrained model.
For CMU-MOSI dataset, we observe 1.69% and 2.44% improvement in accuracy over the Pretrained model performance when only audio and only visual are available, respectively. In all these scenarios, we also see larger improvement in F1 score. The adapted model offers significant improvement when audio-visual modalities are available and text is missing. It shows 6.56% improvement in terms of accuracy and 12.01% improvement in terms of F1 score over the Pretrained model. For CMU-MOSEI dataset, we see even greater improvement in all the metrics. Experiments show 25.7%, 24.21% and 21.41% improvement in accuracy for audio only, visual only and audio-visual scenarios compared to the Pretrained model. We also observe 27.91%-36.30% improvement in F1 score compared to the pretrained model.
We also compare our adaptation method with existing methods for multimodal sentiment analysis. The table shows that for CMU-MOSI dataset, BERT_MAG works better in terms of accuracy but our adaptation method works better in terms of F1 score. One thing to mention is that BERT_MAG uses a pretrained BERT model and finetunes it on the dataset but we are not using any pretraining on extra data. For CMU-MOSEI, our adaptation method works better for most of the cases. Which also demonstrates the generalizability and applicability of our approach to diverse modalities and different tasks.
5 Conclusion
Missing modalities at test time can cause significant degradation in the performance of multimodal systems. In this paper, we presented a simple and parameter-efficient adaptation method for robust multimodal learning with missing modalities. We demonstrated that simple linear operations can efficiently transform a single pretrained multimodal network and achieve performance comparable to multiple (independent) dedicated networks trained for different modality combinations. We evaluated the performance of our method and compared with existing robust methods for different multimodal segmentation and sentiment analysis tasks. Our method requires an extremely small number of additional parameters (e.g., $<0.7\%$ of the total parameters in most experiments), while significantly improving performance compared to existing baseline models and methods for different missing modality scenarios. Our adaptation strategy is applicable to different network architectures and tasks, which can be a versatile solution to build robust multimodal systems. In this work, our main focus was to enhance robustness of existing multimodal models and we did not try to optimize the number of parameters or find the optimal places to insert those learnable layers. Future study will explore these areas to further reduce the number of parameters and application of the approach beyond the tasks and datasets used in this study.
Reproducibility Statement
We are committed to ensuring the reproducibility of our research and to facilitate the broader scientific community in replicating and building upon our work. The source code and trained models are available at github [^1] https://github.com/CSIPlab/Robust-multimodal-learning. We have provided a clear and comprehensive README.md file to guide users in setting up the environment, running the code, and reproducing the results in the paper. We outline the specific data preprocessing steps, list of hyperparameters and configurations used in our experiments in Section 4.2 in the main text and Section 2 in the supplementary section. We hope this makes it easy for others to replicate our experiments. We have provided scripts and instructions in our source code to reproduce the main experimental results presented in this paper. Additionally, we have provided pretrained models allowing others to directly reproduce the results.
Ethics Statement
To the best of our knowledge this work does not give rise to any significant ethical concerns.
Acknowledgements
This work is supported in part by AFOSR award FA9550-21-1-0330 and NSF CAREER award CCF-2046293.
Supplementary Material
1 Datasets
MFNet Dataset introduced by [17], is a popular dataset for RGB-thermal urban scene segmentation, particularly in the context of supporting autonomous driving applications. It comprises a total of 1569 aligned pairs of RGB-thermal images. Within this collection, 820 image pairs were captured during daytime, while 749 pairs were acquired during nighttime. The dataset is divided into distinct training and test sets, each accompanied by pixel-level annotations that define semantic labels for nine classes. Each image is $640\times 480$ pixels.
NYU Depth v2 (NYUDv2) Dataset from [49] is a well-known dataset for RGB-D semantic segmentation. This dataset contains 1449 pairs of aligned RGB-depth images of indoor scenes. The images are divided into training and test sets containing 795 and 654 pairs of images respectively. The dataset also provides per pixel annotations for 13 classes, 40 classes and 894 classes ground truth semantic labels. For our experiments we used the standard 40 classes annotation. Each image is $640\times 480$ pixels and the dataset contains both raw and processed depth maps. For our experiments we used HHA images as proposed by [16] instead of depth maps.
Multimodal Material Segmentation (MCubeS) Dataset was introduced by [32] for accurate multimodal material segmentation with the help of thermal and polarized images alongside RGB images. This dataset has four modalities: RGB, Angle of Linear Polarization, Degree of Linear Polarization and Near-Infrared. Alongside these modalities, the dataset also provides ground truth annotation for semantic and material segmentation. There are 500 image sets divided into train, validation and test sets having 302, 96 and 102 image sets respectively. The images are $1224\times 1024$ pixels each and have 20 classes in total.
CMU-MOSI dataset from [69] is a popularly used for multimodal sentiment analysis. The dataset has 2199 samples each having audio, visual and text as input modalities. It is divided into train, validation and test sets containing 1284, 229 and 686 samples respectively along with annotated sentiment for each sample.
Table 6: Hyperparameters for the experiments on CMU-MOSI and CMU-MOSEI datasets for multimodal sentiment analysis.
Hyperparameters CMU-MOSI CMU-MOSEI
Batch Size 16 4
Initial Learning Rate 0.002 0.0005
Optimizer Adam Adam
Attention Dropout 0.3 0.4
Embedding Dropout 0.2 0.0
Output Dropout 0.5 0.5
Gradient Clip 0.6 0.6
Weight Decay 0.005 0.001
Temporal Conv Kernel Size (T/A/V) 5/5/5 5/1/3
# of Crossmodal Blocks 4 4
CMU-MOSEI is a large scale sentiment analysis dataset from [3]. It is 10 times larger than CMU-MOSI and contains audio, visual and text modalities along with ground truth sentiment annotations. The dataset contains 23453 samples divided into train, validation and test sets for multimodal sentiment analysis and emotion recognition.
2 Implementation Details
We used Python [^2] https://www.python.org/ 3.8.12 and PyTorch [^3] https://pytorch.org/ 1.9.0 to for our implementation. The experiments were done using two NVIDIA RTX 2080 Ti GPUs. We applied automatic mixed precision (AMP) training provided by PyTorch. For CMNeXt model, we use their publicly available code [^4] https://github.com/jamycheung/DELIVER and models trained on all the available modalities for each dataset. We trained the multimodal transformer models on all the modalities using the available code and preprocessed data from the repository [^5] https://github.com/thuiar/MMSA for CMU-MOSI and CMU-MOSEI datasets.
MFNet Dataset: We divied the 4 channel RGB-T images into three channel RGB and one channel thermal images. Then data pre-processing and augmentation was applied following CMNeXt from [70]. MiT-B4 from [62] was the backbone for the base CMNeXt model. One set of scale and shift parameters was learnt for each input modality combination. Input images were sized at $640\times 480$ for both training and testing and we report single scale performance for all the experiments. The scale and shift parameters were trained for 100 epochs with a batch size of 4.
NYUDv2 Dataset: For processing depth maps, we follow SA-Gate by [6] and CMNeXt by [70] and use HHA-encoded images instead of raw depth maps. The already preprocessed dataset can be downloaded from the SA-Gate repository [^6] https://github.com/charlesCXK/RGBD_Semantic_Segmentation_PyTorch. RGB and HHA images were sized at $640\times 480$ pixels each and we used this size for training and testing. The backbone was set to MiT-B4 as suggested in CMNeXt paper. One set of scale and shift parameters was learnt for each input modality combination by feeding available input modalities and setting the missing modality to zero. We train the scale and shift parameters for 100 epochs with a batch size of 4 and report single scale performance.
MCubeS Dataset: We follow the same data pre-processing and augmentations used by the base CMNeXt model from [70]. MiT-B2 from [62] was used as the backbone for this dataset. We set the input image resolution to $512\times 512$ during training and $1024\times 1024$ during testing and report single scale performance with predicted segmentation maps sized at $1024\times 1024$. Similar to other two datasets, we train the learnable parameters for 100 epochs with a batch size of 4.
CMU-MOSI and CMU-MOSEI Datasets: We used Multimodal Transformer (MulT) from [53] as the base model. Preprocessed datasets and all the configurations are available on the repository [^7] https://github.com/thuiar/MMSA. First we trained the multimodal transformer (MulT) model on all the available modalities and then adapted the pretrained model for different missing modality scenarios. The hyperparameters for the experiments are shown in Table 6.
Table 7: Learnable parameter counts for different parameter efficient model adaptation methods. As seen from the table, scale and shift introduce less than 0.7% of the total model parameters.
Method Total Parameters (M) Learnable Parameters (M) % of Total Parameters
Norm 116.560 0.126 0.108
BitFit 116.560 0.378 0.324
LoRA 116.957 0.397 0.340
Scale and Shift 117.349 0.789 0.673
Table 8: Performance comparison with parameter efficient model adaptation techniques on CMNeXt model for MFNet dataset. Average column indicates average performance when one of the two modalities gets missing. Mean accuracy, F1 score and % mIoU are shown for all the experiments.
Methods RGB Thermal Average
mAcc F1 % mIoU mAcc F1 % mIoU mAcc F1 % mIoU
Pretrained 60.74 66.91 53.71 38.18 45.11 35.48 49.46 56.01 44.60
Dedicated 66.28 68.22 55.86 68.35 65.29 53.34 67.32 66.76 54.60
Scale Only 67.09 68.03 54.77 64.00 60.92 49.23 65.55 64.48 52.00
Shift Only 65.82 67.42 54.57 59.77 60.54 48.96 62.80 63.98 51.77
BitFit 66.49 67.40 54.39 61.06 60.59 49.07 63.78 64.00 51.73
LoRA 66.44 67.32 54.19 57.10 59.04 47.45 61.77 63.18 50.82
Norm 66.43 67.07 54.65 57.55 59.22 47.49 61.99 63.15 51.07
Scale and Shift 67.18 68.04 55.22 66.70 62.64 50.89 66.94 65.34 53.06
Table 9: Performance comparison with parameter efficient model adaptation techniques on CMNeXt model for NYUDv2 dataset. Average column indicates average performance when one of the two modalities gets missing. Mean accuracy, F1 score and % mIoU are shown for all the experiments.
Methods RGB Depth Average
mAcc F1 % mIoU mAcc F1 % mIoU mAcc F1 % mIoU
Pretrained 64.10 65.70 51.19 8.30 7.95 5.26 36.20 36.83 28.23
Dedicated 66.00 66.62 52.18 44.80 46.79 33.49 55.40 56.71 42.84
Scale Only 68.18 67.38 53.04 51.54 49.88 36.12 59.86 58.63 44.58
Shift Only 67.54 67.35 53.04 50.30 49.76 36.25 58.92 58.56 44.65
BitFit 67.31 67.33 53.09 50.68 50.27 36.64 59.00 58.80 44.87
LoRA 66.67 67.14 52.87 49.34 48.66 34.97 58.01 57.90 43.92
Norm 67.18 67.34 53.05 48.74 48.06 34.73 57.96 57.70 43.89
Scale and Shift 67.96 67.18 52.82 52.42 50.60 36.72 60.19 58.89 44.77
Table 10: Performance comparison with different parameter efficient model adaptation techniques on CMNeXt model for MCubeS dataset. Average column indicates the average performance. Mean accuracy, F1 score and % mIoU are shown for all the experiments.
Methods RGB RGB-AoLP RGB-AoLP-DoLP Average
mAcc F1 % mIoU mAcc F1 % mIoU mAcc F1 % mIoU mAcc F1 % mIoU
Pretrained 51.63 55.91 42.32 58.66 62.00 48.81 60.06 62.43 49.06 56.78 60.11 46.73
Dedicated 57.70 60.95 48.16 57.56 61.17 48.42 59.12 61.91 49.48 58.13 61.34 48.69
Scale Only 59.64 63.06 50.16 60.28 63.55 50.55 60.96 64.14 51.13 60.29 63.58 50.61
Shift Only 59.82 63.17 50.13 60.10 63.36 50.40 60.61 63.78 50.86 60.18 63.44 50.46
BitFit 59.98 63.24 50.19 60.12 63.52 50.57 60.84 64.03 51.07 60.31 63.60 50.61
LoRA 59.08 62.50 49.59 59.81 63.05 50.07 60.69 63.84 50.80 59.86 63.13 50.15
Norm 59.57 62.89 49.95 60.22 63.49 50.51 60.98 64.08 51.07 60.26 63.49 50.51
Scale and Shift 60.23 63.41 50.43 60.40 63.59 50.62 60.94 64.04 51.11 60.52 63.68 50.72
Table 11: Per class % IoU comparison between pretrained and adapted CMNeXt model on MFNet and MCubeS datasets. Adapted model show better performance for most of the classes leading to overall performance improvement. Here A, D and N stand for Angle of Linear Polarization (AoLP), Degree of Linear Polarization (DoLP) and Near-Infrared (NIR) respectively.
(a) Per class % IoU of pretrained and adapted CMNeXt model on MFNet dataset.
Modalities Methods
Unlabeled
Car
Person
Bike
Curve
Car_Stop
Guardrail
Color_Cone
Bump
Mean
RGB-Thermal CMNeXt 98.31 90.27 74.52 64.52 46.64 39.19 15.09 52.56 59.79 60.10
RGB Pretrained 97.79 87.62 51.13 61.94 30.05 39.36 21.04 45.55 48.95 53.71
Adapted 97.79 88.06 55.55 61.20 34.19 40.52 15.78 48.67 55.21 55.22
Thermal Pretrained 95.97 55.24 68.47 9.27 31.85 2.75 0.0 16.87 38.92 35.48
Adapted 97.46 82.83 70.12 49.03 40.89 26.79 1.84 36.24 52.83 50.89
(b) Per class % IoU of pretrained and adapted CMNeXt model on MCubeS dataset.
Modalities Methods
Asphalt
Concrete
Metal
Road_Marking
Fabric
Glass
Plaster
Plastic
Rubber
Sand
Gravel
Ceramic
Cobblestone
Brick
Grass
Wood
Leaf
Water
Human
Sky
Mean
RGB-A-D-N CMNeXt 84.4 44.9 53.9 74.6 32.1 54.0 0.8 28.7 29.8 67.0 66.2 27.7 68.5 42.8 58.7 49.7 75.3 55.6 19.1 96.52 51.5
RGB Pretrained 69.7 39.2 47.6 67.3 26.9 44.6 0.2 20.9 15.2 61.8 36.7 19.1 67.2 36.0 49.5 36.1 71.6 36.1 14.7 86.3 42.3
Adapted 85.8 43.7 52.6 73.8 27.9 51.0 0.8 24.2 30.4 67.8 72.9 27.1 68.1 42.9 57.6 49.0 74.9 43.4 18.3 96.5 50.4
RGB-A Pretrained 83.2 43.3 50.7 72.6 26.4 51.9 0.2 28.1 22.2 67.7 63.4 22.7 67.5 40.6 54.4 44.9 73.9 44.8 21.8 96.0 48.8
Adapted 84.4 45.4 53.8 74.5 30.4 53.2 0.6 26.9 28.8 69.0 69.3 24.8 67.5 43.2 58.4 48.2 75.1 48.1 14.4 96.4 50.6
RGB-A-D Pretrained 84.5 41.2 46.7 72.8 25.2 51.6 0.3 26.1 28.8 66.7 65.6 26.0 66.5 40.4 50.0 45.1 72.7 49.4 25.6 96.3 49.1
Adapted 84.1 45.6 54.1 74.6 30.5 54.2 0.6 28.1 30.1 69.0 67.6 25.9 67.8 43.8 58.0 49.1 75.0 53.7 13.7 96.5 51.1
3 Number of Learnable Parameters
We report the number of learnable parameters for different parameter-efficient adaptation techniques (for multimodal segmentation) in Table 7. We insert scale and shift layers after each linear, convolutional and norm (both batch norm and layer norm) layers. The number of learnable parameter varies with the size of the backbone. We used MiT-B4 as the backbone while counting these learnable parameters. Scale and shift adds only 0.789M learnable parameters which is less than 0.7% of the total model parameters. Despite this very few parameters, it improves performance significantly in different missing modality scenarios. For this study we mainly focused on improving missing modality robustness and did not try to optimize the number of learnable parameters. We will leave that part for future studies.
Figure 3: Visualization of predicted segmentation maps for pretrained and adapted models on MFNet and NYUDv2 datasets for multimodal semantic segmentation and MCubeS dataset for multimodal material segmentation. Only RGB images are shown from MCubeS dataset for brevity. CMNeXt column shows the predictions when all the modalities are available. Segmentation quality improves significantly after model adaptation for all the input modality combinations. A, D and N stand for angle of linear polarization, degree of linear polarization and near-infrared respectively.
4 Performance Comparison with Parameter Efficient Model Adaption Techniques
We performed a detailed performance comparison with other parameter efficient methods for the three segmentation datasets. The results are summarized in Table 8 for RGB-thermal segmentation on MFNet dataset, Table 9 for RGB-depth segmentation on NYUDv2 dataset and Table 10 for multimodal material segmentation on MCubeS dataset. For each method, we take a model trained on all the available modalities. Then we freeze the pretrained weights and tune the learnable parameters for the corresponding adaption method. We have shown mean accuracy, F1 score and % mIoU for each experiment.
4.1 Performance Comparison for RGB-Thermal Semantic Segmentation on MFNet Dataset
Table 8 summarizes the results on MFNet dataset when the base CMNeXt model is adapted with other parameter efficient model adaptation techniques. Experiments show that scale and shift shows the best performance in all three matrices compared to all other methods. It shows a significant improvement of +8.46% in mIoU, +9.33% in F1 score and +17.48% in mean accuracy on an average over the pretrained model. The average performance is also close to dedicatedly trained models.
4.2 Performance Comparison for RGB-Depth Semantic Segmentation on NYUDv2 Dataset
Similar trend is observed for RGB-Depth semantic segmentation on NYUDv2 dataset as shown in Table 9. Scale only and BitFit adapted models show slightly better performance for some of the matrices. But in most of the cases scale and shift adapted model performs better. For all the matrices, scale and shift shows a significant improvement of +16.54% in mIoU, +22.05% in F1 score and +23.99% in mean accuracy over the pretrained model on an average and consistently outperforms dedicated training.
4.3 Performance Comparison for Multimodal Material Segmentation on MCubeS Dataset
We show comparison with parameter efficient model adaptation techniques on MCubeS dataset in Table 10. Scale and shift outperforms all other methods in most of the matrices for all input combinations. It also shows an improvement of +3.99% in mIoU, +3.57% in F1 score and +3.74% in mean accuracy on an average over the pretrained model. Furthermore, Scale and shift also outperforms dedicated training for all input modality combinations. These experiments corroborate the fact that scale and shift provides better model adaption for different missing modality scenarios.
5 Per Class IoU Comparison
To further analyze how the adaption is helping the model improve overall semantic and material segmentation performance, we conduct a per-class % intersection over union (IoU) analysis on the pretrained and adapted models. Table 11(a) and 11(b) summarize the results. We show the per class % IoU comparison for different missing modality situations on MFNet dataset on Table 11(a). From the table we can see that when RGB is available and thermal is missing, the adaptation helps improve performance for most of the classes. Though we see some performance drop for bike (-0.74%) and guardrail (-5.25%) classes, the rest of the classes have better % IoU than the pretrained model. Bump (+6.26%), person (+4.42%), and curve (+4.14%) classes show greater improvement after adaptation. When thermal is available and RGB is missing, adaptation improves performance for all the classes. Among the classes, bike (+39.76%), car (+27.59%), car stop (+24.04%), color cone (+19.37%) and bump (+13.91%) are showing impressive performance improvement over the pretrained model.
Results for MCubeS dataset is shown on Table 11(b). Here A, D, and N stand for angle of linear polarization (AoLP), degree of linear polarization (DoLP) and near-infrared (NIR) respectively. Experiments show that when only RGB is available and the rest of the modalities are missing, the adapted model performs better in detecting all the 20 classes present in the dataset. Gravel (36.2%), asphalt (16.1%), rubber (15.2%), wood (12.9%) and sky (10.2%) are some of the classes who show the most performance boost after adaptation. In other input combinations, most of the classes see performance improvement compared to the pretrained model. Though we see some performance drop in a few classes, most of the classes show improvement in % IoU which leads to the overall performance improvement after adaption.
6 Visualization of Predicted Segmentation Maps
We show the predicted segmentation maps from the pretrained and adapted models in Figure 3. For each dataset, we show the input images, predictions from the base CMNeXt model when all the modalities are available, predictions from the adapted and pretrained models for different missing modality scenarios. For brevity, we only show RGB input images for MCubeS dataset. A, D and N stand for angle of linear polarization (AoLP), degree of linear polarization (DoLP) and near-infrared (NIR) respectively. Modalities that are available during testing are shown in parenthesis while other modalities are missing.
For MFNet dataset, Figure LABEL:fig:vis-mfnet-2 shows that when only RGB is available, the pretrained model performs very poorly in detecting humans. On the other hand, if only thermal is available, the pretrained model can not detect cars very accurately. But the adapted model can detect both humans and cars more accurately in both of the scenarios. In all the cases, the predictions form the adapted model is closer to the predictions of the base CMNeXt model when all the modalities are available.
Predictions from NYUDv2 dataset is shown on Figure LABEL:fig:vis-nyu-2. We can see that the adapted model can identify bed, furniture and other classes more accurately than the pretrained model for different missing modality scenarios. The pretrained model performs very poorly when only depth is available and RGB is missing. But detection accuracy improves significantly after model adaptation. For MCubeS dataset, as seen in Figure LABEL:fig:vis-mcubes-2, predictions from the pretrained model shows artifacts when detecting different materials. On the other hand, the adapted model is showing more accuracy in detecting sky, cobblestone, sand and brick. For all the three datasets, the predictions from the adapted model is more accurate and closer to the all modality predictions of the base CMNeXt model.